OpenStudy (anonymous):

Hannah has deposited $742 in a savings account that earns interest at a rate of 3.9% compounded twice a year. What will the account balance be in 11 years?

OpenStudy (anonymous):

B = D (1 + r/n)^nt = 742 ( 1+ .039/2)^2*11 = $1134.81

OpenStudy (anonymous):

thank you so much! can you help with my other ones please?